OpenStudy (anonymous):

I need the answer to a algebra problem. lnx-ln3=18

OpenStudy (anonymous):

lnx=ln3+18. Raise e to both sides. e^lnx=(e^ln3)(e^18) So, x=3e^18

OpenStudy (anonymous):

Thank You